# Jacobian Descent For Multi-Objective Optimization

### v1

@@ A.3 Theorem 1
We first bound, for any $t\vargeq 0$, ${\boldsymbol{1}}^{\top}({\boldsymbol{f}}({\boldsymbol{x}}_{t+1})-{\boldsymbol{f}}({\boldsymbol{x}}_{t}))$ as follows
${\boldsymbol{1}}^{\top}({\boldsymbol{f}}({\boldsymbol{x}}_{t+1})-{\boldsymbol{f}}({\boldsymbol{x}}_{t}))$
@@
Summing this over $t=0,\dots,T-1$ yields
$\frac{1}{2\beta\sqrt{m}}\sum_{t=0}^{T-1}{\boldsymbol{w}}_{t}^{\top}G_{t}{\boldsymbol{w}}_{t}$
$\sum_{t=0}^{T-1}{\boldsymbol{1}}^{\top}({\boldsymbol{f}}({\boldsymbol{x}}_{t})-{\boldsymbol{f}}({\boldsymbol{x}}_{t+1}))$
@@
${\boldsymbol{1}}^{\top}({\boldsymbol{f}}({\boldsymbol{x}}_{0})-{\boldsymbol{f}}({\boldsymbol{x}}^{\ast})).$ $\left(\begin{subarray}{c}\text{Assumption}\\ {\boldsymbol{1}}^{\top}{\boldsymbol{f}}({\boldsymbol{x}}^{\ast})\varleq{\boldsymbol{1}}^{\top}{\boldsymbol{f}}({\boldsymbol{x}}_{T})\end{subarray}\right)$ (17)
Since ${\boldsymbol{0}}\preceq{\boldsymbol{w}}_{t}$,
${\boldsymbol{w}}_{t}^{\top}({\boldsymbol{f}}({\boldsymbol{x}}_{t})-{\boldsymbol{f}}({\boldsymbol{x}}^{\ast}))$
@@
Summing this over $t=0,\dots,T-1$ yields
$\sum_{t=0}^{T-1}{\boldsymbol{w}}_{t}^{\top}({\boldsymbol{f}}({\boldsymbol{x}}_{t})-{\boldsymbol{f}}({\boldsymbol{x}}^{\ast}))$
$\frac{1}{2\beta\sqrt{m}}\sum_{t=0}^{\top}{\boldsymbol{w}}_{t}^{\top}G_{t}{\boldsymbol{w}}_{t}+\frac{\beta\sqrt{m}}{2}\bigg{(}\|{\boldsymbol{x}}_{0}-{\boldsymbol{x}}^{\ast}\|^{2}-\|{\boldsymbol{x}}_{T}-{\boldsymbol{x}}^{\ast}\|^{2}\bigg{)}$ (Telescoping $\left(\begin{subarray}{c}\text{Telescoping sum}\end{subarray}\right)$
$\frac{1}{2\beta\sqrt{m}}\sum_{t=0}^{\top}{\boldsymbol{w}}_{t}^{\top}G_{t}{\boldsymbol{w}}_{t}+\frac{\beta\sqrt{m}}{2}\|{\boldsymbol{x}}_{0}-{\boldsymbol{x}}^{\ast}\|^{2}$
@@
$\frac{1}{T}\sum_{t=0}^{T-1}{\boldsymbol{w}}_{t}^{\top}({\boldsymbol{f}}({\boldsymbol{x}}_{t})-{\boldsymbol{f}}({\boldsymbol{x}}^{\ast}))\varleq\frac{1}{T}\bigg{(}{\boldsymbol{1}}^{\top}({\boldsymbol{f}}({\boldsymbol{x}}_{0})-{\boldsymbol{f}}({\boldsymbol{x}}^{\ast}))+\frac{\beta\sqrt{m}}{2}\|{\boldsymbol{x}}_{0}-{\boldsymbol{x}}^{\ast}\|^{2}\bigg{)},$
@@
For all $t\vargeq 0$, let $J_{t}=\mathcal{J}\kern-2.0pt{\boldsymbol{f}}({\boldsymbol{x}}_{t})$, $G_{t}=J_{t}J_{t}^{\top}$. Then
${\boldsymbol{x}}_{t+1}$ $={\boldsymbol{x}}_{t}-\eta\mathcal{A}_{\text{UPGrad}}(J_{t})$
$={\boldsymbol{x}}_{t}-\eta J_{t}^{\top}{\boldsymbol{w}}_{t}.$
Substituting ${\boldsymbol{w}}={\boldsymbol{1}}$ in the term $\frac{{\boldsymbol{1}}}{2\sqrt{m}}-\frac{{\boldsymbol{w}}}{\|{\boldsymbol{w}}\|}$ of Lemma 6 yields
$\frac{{\boldsymbol{1}}}{2\sqrt{m}}-\frac{{\boldsymbol{w}}}{\|{\boldsymbol{w}}\|}$ $=-\frac{{\boldsymbol{1}}}{2\sqrt{m}}$
$\prec{\boldsymbol{0}}.$
Therefore there exists some $\varepsilon>0$ such that any ${\boldsymbol{w}}\in{\mathbb{R}}^{m}$ with $\|{\boldsymbol{1}}-{\boldsymbol{w}}\|<\varepsilon$ satisfies $\frac{{\boldsymbol{1}}}{2\sqrt{m}}\prec\frac{{\boldsymbol{w}}}{\|{\boldsymbol{w}}\|}$. Denote by $B_{\varepsilon}({\boldsymbol{1}})=\{{\boldsymbol{w}}\in{\mathbb{R}}^{m}:\|{\boldsymbol{1}}-{\boldsymbol{w}}\|<\varepsilon\}$, i.e. for all ${\boldsymbol{w}}\in B_{\varepsilon}({\boldsymbol{1}})$, $\frac{{\boldsymbol{1}}}{2\sqrt{m}}\prec\frac{{\boldsymbol{w}}}{\|{\boldsymbol{w}}\|}$. By the non-conflicting property of $\mathcal{A}_{\text{UPGrad}}$, ${\boldsymbol{0}}\preceq G_{t}{\boldsymbol{w}}_{t}$ and therefore for all ${\boldsymbol{w}}\in B_{\varepsilon}({\boldsymbol{1}})$,
${\boldsymbol{w}}^{\top}({\boldsymbol{f}}({\boldsymbol{x}}_{t+1})-{\boldsymbol{f}}({\boldsymbol{x}}_{t}))$
@@
Since ${\boldsymbol{w}}^{\top}{\boldsymbol{f}}({\boldsymbol{x}}_{t})$ is bounded and non-increasing, it converges. Since $B_{\varepsilon}({\boldsymbol{1}})$ contains a basis of ${\mathbb{R}}^{m}$, ${\boldsymbol{f}}({\boldsymbol{x}}_{t})$ converges to some ${\boldsymbol{f}}^{\ast}\in{\mathbb{R}}^{m}$. By assumption on ${\boldsymbol{f}}$, there exists ${\boldsymbol{x}}^{\ast}$ in the Pareto set such that ${\boldsymbol{f}}({\boldsymbol{x}}^{\ast})\preceq{\boldsymbol{f}}^{\ast}$.
We now prove that ${\boldsymbol{f}}({\boldsymbol{x}}^{\ast})={\boldsymbol{f}}^{\ast}$. Since ${\boldsymbol{f}}({\boldsymbol{x}}^{\ast})\preceq{\boldsymbol{f}}^{\ast}$, it is sufficient to show that ${\boldsymbol{1}}^{\top}({\boldsymbol{f}}^{\ast}-{\boldsymbol{f}}({\boldsymbol{x}}^{\ast}))\varleq 0$.
First, the additional assumption of Lemma 7 applies since ${\boldsymbol{1}}^{\top}{\boldsymbol{f}}({\boldsymbol{x}}_{t})$ decreases to ${\boldsymbol{1}}^{\top}{\boldsymbol{f}}^{\ast}$ which is larger than ${\boldsymbol{1}}^{\top}{\boldsymbol{f}}({\boldsymbol{x}}^{\ast})$. Therefore
${\boldsymbol{1}}^{\top}({\boldsymbol{f}}^{\ast}-{\boldsymbol{f}}({\boldsymbol{x}}^{\ast}))$
@@
${\boldsymbol{1}}^{\top}({\boldsymbol{f}}^{\ast}-{\boldsymbol{f}}({\boldsymbol{x}}^{\ast}))$
$\lim_{T\to\infty}\frac{1}{T}\sum_{t=0}^{T-1}{\boldsymbol{w}}_{t}^{\top}({\boldsymbol{f}}^{\ast}-{\boldsymbol{f}}({\boldsymbol{x}}_{t}))$
$\lim_{T\to\infty}\frac{1}{T}\sum_{t=0}^{T-1}\|{\boldsymbol{w}}_{t}\|\cdot\|{\boldsymbol{f}}^{\ast}-{\boldsymbol{f}}({\boldsymbol{x}}_{t})\|$ $\left(\begin{subarray}{c}\text{Cauchy-Schwartz}\\ \text{inequality}\end{subarray}\right)$
@@ Appendix B Properties of existing aggregators
In the following, we prove the properties of the aggregators from Table 1. Some aggregators, e.g. $\mathcal{A}_{\text{RGW}}$, $\mathcal{A}_{\text{GradDrop}}$ and $\mathcal{A}_{\text{PCGrad}}$, are non-deterministic and are thus not technically functions but rather random variables whose distribution depends on the matrix $J\in\mathbb{R}^{m\times n}$ to aggregate. Still, the properties of Section 2.2 can be easily adapted to a random setting. If $\mathcal{A}$ is a random aggregator, then for any $J$, $\mathcal{A}(J)$ is a random vector in ${\mathbb{R}}^{n}$. The aggregator is non-conflicting if $\mathcal{A}(J)$ is in the dual cone of the rows of $J$ with probability 1. It is linear under scaling if for all $J\in\mathbb{R}^{m\times n}$, there is a – possibly random – matrix $\mathfrak{J}\in\mathbb{R}^{m\times n}$, such that for all ${\boldsymbol{0}}\prec{\boldsymbol{c}}\in{\mathbb{R}}^{m}$, $\mathcal{A}(\operatorname*{diag}({\boldsymbol{c}})\cdot J)=\mathfrak{J}^{\top}\cdot{\boldsymbol{c}}$. Finally, $\mathcal{A}$ is weighted if for any $J\in\mathbb{R}^{m\times n}$ there is, with probability 1, some weighting ${\boldsymbol{w}}\in{\mathbb{R}}^{m}$ such that $\mathcal{A}(J)=J^{\top}\cdot{\boldsymbol{w}}$.
@@ B.1 Mean
$\mathcal{A}_{\text{Mean}}$ simply averages the rows of the input matrix, i.e. for all $J\in\mathbb{R}^{m\times n}$, $\mathcal{A}_{\text{Mean}}(J)=\frac{1}{m}J^{\top}\cdot{\boldsymbol{1}}$. It is weighted with constant weighting equal to $\frac{1}{m}{\boldsymbol{1}}$. For any ${\boldsymbol{c}}\in{\mathbb{R}}^{m}$, $\mathcal{A}_{\text{Mean}}(\operatorname*{diag}({\boldsymbol{c}})\cdot J)=J^{\top}\cdot{\boldsymbol{c}}$, which is linear in ${\boldsymbol{c}}$. $\mathcal{A}_{\text{Mean}}$ is therefore linear under scaling. $\mathcal{A}_{\text{Mean}}\left(\begin{bmatrix}-2\\ 4\end{bmatrix}\right)=\begin{bmatrix}1\end{bmatrix}$, which conflicts with $\begin{bmatrix}-2\end{bmatrix}$, so $\mathcal{A}_{\text{Mean}}$ is not non-conflicting.
@@ B.2 MGDA
The optimization algorithm presented in [6], called MGDA, is tied to a particular method for aggregating the gradients. We thus refer to this aggregator as $\mathcal{A}_{\text{MGDA}}$. The dual problem of this method was also introduced independently in [5]. We show the equivalence between the two solutions to make the analysis of $\mathcal{A}_{\text{MGDA}}$ easier.
Let $J\in\mathbb{R}^{m\times n}$. The aggregation described in [6] is defined by the weighting that is a solution to
$\operatorname*{arg\,min}_{\begin{subarray}{c}{\boldsymbol{0}}\preceq{\boldsymbol{w}}:\\ {\boldsymbol{1}}^{\top}{\boldsymbol{w}}=1\end{subarray}}\|J^{\top}{\boldsymbol{w}}\|^{2}$ (18)
By construction, $\mathcal{A}_{\text{MGDA}}$ is thus weighted.
In Equation (3) of [5], the following problem is studied
$\min_{\begin{subarray}{c}\alpha\in\mathbb{R},{\boldsymbol{x}}\in{\mathbb{R}}^{n}:\\ J{\boldsymbol{x}}\preceq\alpha{\boldsymbol{1}}\end{subarray}}\alpha+\frac{1}{2}\|{\boldsymbol{x}}\|^{2}$ (19)
We show that the problems in (18) and (19) are dual to each other. Furthermore, the duality gap is null since this is a convex problem.
The Lagrangian of the problem in (19) is given by $\mathcal{L}(\alpha,{\boldsymbol{x}},{\boldsymbol{\mu}})=\alpha+\frac{1}{2}\|{\boldsymbol{x}}\|^{2}-{\boldsymbol{\mu}}^{\top}(\alpha{\boldsymbol{1}}-J{\boldsymbol{x}})$. Differentiating w.r.t. $\alpha$ and ${\boldsymbol{x}}$ gives respectively $1-{\boldsymbol{\mu}}^{\top}{\boldsymbol{1}}$ and ${\boldsymbol{x}}+J^{\top}{\boldsymbol{\mu}}$. The dual problem is obtained by setting those two to ${\boldsymbol{0}}$ and then maximizing the Lagrangian on ${\boldsymbol{0}}\preceq{\boldsymbol{\mu}}$ and $\alpha$, i.e.
$\operatorname*{arg\,max}_{\begin{subarray}{c}\alpha,{\boldsymbol{0}}\preceq{\boldsymbol{\mu}}:\\ {\boldsymbol{\mu}}^{\top}{\boldsymbol{1}}=1\end{subarray}}\alpha+\frac{1}{2}\|J^{\top}{\boldsymbol{\mu}}\|^{2}-{\boldsymbol{\mu}}^{\top}(\alpha{\boldsymbol{1}}+JJ^{\top}{\boldsymbol{\mu}})$
@@
Therefore, (18) and (19) are equivalent, with ${\boldsymbol{x}}=-J^{\top}{\boldsymbol{w}}$.
Observe that since in (19), $\alpha=0$ and ${\boldsymbol{x}}={\boldsymbol{0}}$ is feasible, the objective is non-positive and therefore $\alpha\varleq 0$. Substituting ${\boldsymbol{x}}=-J^{\top}{\boldsymbol{w}}$ in $J\cdot{\boldsymbol{x}}\preceq\alpha{\boldsymbol{1}}\preceq{\boldsymbol{0}}$ yields ${\boldsymbol{0}}\preceq JJ^{\top}{\boldsymbol{w}}$, i.e. ${\boldsymbol{0}}\preceq J\cdot\mathcal{A}_{\text{MGDA}}(J)$, so $\mathcal{A}_{\text{MGDA}}$ is non-conflicting.
With $J=\begin{bmatrix}2&0\\ 0&2\\ a&a\end{bmatrix}$, if $0\varleq a\varleq 1$, $\mathcal{A}_{\text{MGDA}}(J)=\begin{bmatrix}a\\ a\end{bmatrix}$. However, if $a\vargeq 1$, $\mathcal{A}_{\text{MGDA}}(J)=\begin{bmatrix}1\\ 1\end{bmatrix}$. This is not affine in $a$, so $\mathcal{A}_{\text{MGDA}}$ is not linear under scaling. In particular, if any row of $J$ is ${\boldsymbol{0}}$, $\mathcal{A}_{\text{MGDA}}(J)={\boldsymbol{0}}$. This implies that the optimization will stop whenever one objective has converged.

### v2

@@ A.3 Theorem 1
We first bound, for any $t\in\mathbb{N}$, ${\bm{1}}^{\top}\big{(}{\bm{f}}({\bm{x}}_{t+1})-{\bm{f}}({\bm{x}}_{t})\big{)}$ as follows
${\bm{1}}^{\top}\big{(}{\bm{f}}({\bm{x}}_{t+1})-{\bm{f}}({\bm{x}}_{t})\big{)}$
@@
Summing this over $t\in[T]$ yields
$\frac{1}{2\beta\sqrt{m}}\sum_{t\in[T]}{\bm{w}}_{t}^{\top}G_{t}{\bm{w}}_{t}$
$\sum_{t\in[T]}{\bm{1}}^{\top}\big{(}{\bm{f}}({\bm{x}}_{t})-{\bm{f}}({\bm{x}}_{t+1})\big{)}$
@@
${\bm{1}}^{\top}\big{(}{\bm{f}}({\bm{x}}_{1})-{\bm{f}}({\bm{x}}^{\ast})\big{)}.$ $\left(\begin{subarray}{c}\text{Assumption}\\ {\bm{1}}^{\top}{\bm{f}}({\bm{x}}^{\ast})\varleq{\bm{1}}^{\top}{\bm{f}}({\bm{x}}_{T+1})\end{subarray}\right)$ (17)
Since ${\bm{0}}\preceq{\bm{w}}_{t}$,
${\bm{w}}_{t}^{\top}\big{(}{\bm{f}}({\bm{x}}_{t})-{\bm{f}}({\bm{x}}^{\ast})\big{)}$
@@
Summing this over $t\in[T]$ yields
$\sum_{t\in[T]}{\bm{w}}_{t}^{\top}\big{(}{\bm{f}}({\bm{x}}_{t})-{\bm{f}}({\bm{x}}^{\ast})\big{)}$
$\frac{1}{2\beta\sqrt{m}}\sum_{t\in[T]}{\bm{w}}_{t}^{\top}G_{t}{\bm{w}}_{t}+\frac{\beta\sqrt{m}}{2}\Big{(}\|{\bm{x}}_{1}-{\bm{x}}^{\ast}\|^{2}-\|{\bm{x}}_{T+1}-{\bm{x}}^{\ast}\|^{2}\Big{)}$ (Telescoping $\left(\begin{subarray}{c}\text{Telescoping sum}\end{subarray}\right)$
$\frac{1}{2\beta\sqrt{m}}\sum_{t\in[T]}{\bm{w}}_{t}^{\top}G_{t}{\bm{w}}_{t}+\frac{\beta\sqrt{m}}{2}\|{\bm{x}}_{1}-{\bm{x}}^{\ast}\|^{2}$
@@
$\frac{1}{T}\sum_{t\in[T]}{\bm{w}}_{t}^{\top}\big{(}{\bm{f}}({\bm{x}}_{t})-{\bm{f}}({\bm{x}}^{\ast})\big{)}\varleq\frac{1}{T}\bigg{(}{\bm{1}}^{\top}({\bm{f}}({\bm{x}}_{1})-{\bm{f}}({\bm{x}}^{\ast}))+\frac{\beta\sqrt{m}}{2}\|{\bm{x}}_{1}-{\bm{x}}^{\ast}\|^{2}\bigg{)},$
@@
For all $t\in\mathbb{N}$, let $J_{t}=\mathcal{J}\kern-2.0pt{\bm{f}}({\bm{x}}_{t})$, $G_{t}=J_{t}J_{t}^{\top}$. Then
${\bm{x}}_{t+1}$ $={\bm{x}}_{t}-\eta\mathcal{A}_{\text{UPGrad}}(J_{t})$
$={\bm{x}}_{t}-\eta J_{t}^{\top}{\bm{w}}_{t}.$
Substituting ${\bm{w}}={\bm{1}}$ in the term $\frac{{\bm{1}}}{2\sqrt{m}}-\frac{{\bm{w}}}{\|{\bm{w}}\|}$ of Lemma 6 yields
$\frac{{\bm{1}}}{2\sqrt{m}}-\frac{{\bm{w}}}{\|{\bm{w}}\|}$ $=-\frac{{\bm{1}}}{2\sqrt{m}}$
$\prec{\bm{0}}.$
Therefore there exists some $\varepsilon>0$ such that any ${\bm{w}}\in{\mathbb{R}}^{m}$ with $\|{\bm{1}}-{\bm{w}}\|<\varepsilon$ satisfies $\frac{{\bm{1}}}{2\sqrt{m}}\prec\frac{{\bm{w}}}{\|{\bm{w}}\|}$. Denote by $B_{\varepsilon}({\bm{1}})=\{{\bm{w}}\in{\mathbb{R}}^{m}:\|{\bm{1}}-{\bm{w}}\|<\varepsilon\}$, i.e. for all ${\bm{w}}\in B_{\varepsilon}({\bm{1}})$, $\frac{{\bm{1}}}{2\sqrt{m}}\prec\frac{{\bm{w}}}{\|{\bm{w}}\|}$. By the non-conflicting property of $\mathcal{A}_{\text{UPGrad}}$, ${\bm{0}}\preceq G_{t}{\bm{w}}_{t}$ and therefore for all ${\bm{w}}\in B_{\varepsilon}({\bm{1}})$,
${\bm{w}}^{\top}\big{(}{\bm{f}}({\bm{x}}_{t+1})-{\bm{f}}({\bm{x}}_{t})\big{)}$
@@
Since ${\bm{w}}^{\top}{\bm{f}}({\bm{x}}_{t})$ is bounded and non-increasing, it converges. Since $B_{\varepsilon}({\bm{1}})$ contains a basis of ${\mathbb{R}}^{m}$, ${\bm{f}}({\bm{x}}_{t})$ converges to some ${\bm{f}}^{\ast}\in{\mathbb{R}}^{m}$. By assumption on ${\bm{f}}$, there exists ${\bm{x}}^{\ast}$ in the Pareto set satisfying ${\bm{f}}({\bm{x}}^{\ast})\preceq{\bm{f}}^{\ast}$.
We now prove that ${\bm{f}}({\bm{x}}^{\ast})={\bm{f}}^{\ast}$. Since ${\bm{f}}({\bm{x}}^{\ast})\preceq{\bm{f}}^{\ast}$, it is sufficient to show that ${\bm{1}}^{\top}\big{(}{\bm{f}}^{\ast}-{\bm{f}}({\bm{x}}^{\ast})\big{)}\varleq 0$.
First, the additional assumption of Lemma 7 applies since ${\bm{1}}^{\top}{\bm{f}}({\bm{x}}_{t})$ decreases to ${\bm{1}}^{\top}{\bm{f}}^{\ast}$ which is larger than ${\bm{1}}^{\top}{\bm{f}}({\bm{x}}^{\ast})$. Therefore
${\bm{1}}^{\top}\big{(}{\bm{f}}^{\ast}-{\bm{f}}({\bm{x}}^{\ast})\big{)}$
@@
${\bm{1}}^{\top}\big{(}{\bm{f}}^{\ast}-{\bm{f}}({\bm{x}}^{\ast})\big{)}$
$\lim_{T\to\infty}\frac{m}{T}\sum_{t\in[T]}{\bm{w}}_{t}^{\top}\big{(}{\bm{f}}^{\ast}-{\bm{f}}({\bm{x}}_{t})\big{)}$
$\lim_{T\to\infty}\frac{m}{T}\sum_{t\in[T]}\|{\bm{w}}_{t}\|\cdot\big{\|}{\bm{f}}^{\ast}-{\bm{f}}({\bm{x}}_{t})\big{\|}$ $\left(\begin{subarray}{c}\text{Cauchy-Schwartz}\\ \text{inequality}\end{subarray}\right)$
@@ Appendix B Properties of existing aggregators
In the following, we prove the properties of the aggregators from Table 1. Some aggregators, e.g. $\mathcal{A}_{\text{RGW}}$, $\mathcal{A}_{\text{GradDrop}}$ and $\mathcal{A}_{\text{PCGrad}}$, are non-deterministic and are thus not technically functions but rather random variables whose distribution depends on the matrix $J\in\mathbb{R}^{m\times n}$ to aggregate. Still, the properties of Section 2.2 can be easily adapted to a random setting. If $\mathcal{A}$ is a random aggregator, then for any $J$, $\mathcal{A}(J)$ is a random vector in ${\mathbb{R}}^{n}$. The aggregator is non-conflicting if $\mathcal{A}(J)$ is in the dual cone of the rows of $J$ with probability 1. It is linear under scaling if for all $J\in\mathbb{R}^{m\times n}$, there is a – possibly random – matrix $\mathfrak{J}\in\mathbb{R}^{m\times n}$ such that for all ${\bm{0}}\prec{\bm{c}}\in{\mathbb{R}}^{m}$, $\mathcal{A}(\operatorname*{diag}({\bm{c}})\cdot J)=\mathfrak{J}^{\top}\cdot{\bm{c}}$. Finally, $\mathcal{A}$ is weighted if for any $J\in\mathbb{R}^{m\times n}$ there is a – possibly random – weighting ${\bm{w}}\in{\mathbb{R}}^{m}$ satisfying $\mathcal{A}(J)=J^{\top}\cdot{\bm{w}}$.
@@ B.1 Mean
$\mathcal{A}_{\text{Mean}}$ simply averages the rows of the input matrix, i.e. for all $J\in\mathbb{R}^{m\times n}$,
$\mathcal{A}_{\text{Mean}}(J)=\frac{1}{m}J^{\top}\cdot{\bm{1}}$ (18)
  Non-conflicting.
$\mathcal{A}_{\text{Mean}}\left(\begin{bmatrix}-2\\ 4\end{bmatrix}\right)=\begin{bmatrix}1\end{bmatrix}$, which conflicts with $\begin{bmatrix}-2\end{bmatrix}$, so $\mathcal{A}_{\text{Mean}}$ is not non-conflicting.
  Linear under scaling.
For any ${\bm{c}}\in{\mathbb{R}}^{m}$, $\mathcal{A}_{\text{Mean}}(\operatorname*{diag}({\bm{c}})\cdot J)=\frac{1}{m}J^{\top}\cdot{\bm{c}}$, which is linear in ${\bm{c}}$. $\mathcal{A}_{\text{Mean}}$ is therefore linear under scaling.
  Weighted.
By (18), $\mathcal{A}_{\text{Mean}}$ is weighted with constant weighting equal to $\frac{1}{m}{\bm{1}}$.
@@ B.2 MGDA
The optimization algorithm presented in Désidéri (2012), called MGDA, is tied to a particular method for aggregating the gradients. We thus refer to this aggregator as $\mathcal{A}_{\text{MGDA}}$. The dual problem of this method was also introduced independently in Fliege & Svaiter (2000). We show the equivalence between the two solutions to make the analysis of $\mathcal{A}_{\text{MGDA}}$ easier.
For all $J\in\mathbb{R}^{m\times n}$, the aggregation described in Désidéri (2012) is defined as
$\mathcal{A}_{\text{MGDA}}(J)$ $=J^{\top}\cdot{\bm{w}}$ (19)
$\text{with}\hskip 12.80365pt{\bm{w}}$ $\in\operatorname*{arg\,min}_{\begin{subarray}{c}{\bm{0}}\preceq{\bm{v}}:\\ {\bm{1}}^{\top}{\bm{v}}=1\end{subarray}}\left\|J^{\top}{\bm{v}}\right\|^{2}$ (20)
In Equation (3) of Fliege & Svaiter (2000), the following problem is studied:
$\min_{\begin{subarray}{c}\alpha\in\mathbb{R},{\bm{x}}\in{\mathbb{R}}^{n}:\\ J{\bm{x}}\preceq\alpha{\bm{1}}\end{subarray}}\alpha+\frac{1}{2}\|{\bm{x}}\|^{2}$ (21)
We show that the problems in (20) and (21) are dual to each other. Furthermore, the duality gap is null since this is a convex problem. The Lagrangian of the problem in (21) is given by $\mathcal{L}(\alpha,{\bm{x}},{\bm{\mu}})=\alpha+\frac{1}{2}\|{\bm{x}}\|^{2}-{\bm{\mu}}^{\top}(\alpha{\bm{1}}-J{\bm{x}})$. Differentiating w.r.t. $\alpha$ and ${\bm{x}}$ gives respectively $1-{\bm{1}}^{\top}{\bm{\mu}}$ and ${\bm{x}}+J^{\top}{\bm{\mu}}$. The dual problem is obtained by setting those two to ${\bm{0}}$ and then maximizing the Lagrangian on ${\bm{0}}\preceq{\bm{\mu}}$ and $\alpha$, i.e.
$\operatorname*{arg\,max}_{\begin{subarray}{c}\alpha,{\bm{0}}\preceq{\bm{\mu}}:\\ {\bm{1}}^{\top}{\bm{\mu}}=1\end{subarray}}\alpha+\frac{1}{2}\left\|J^{\top}{\bm{\mu}}\right\|^{2}-{\bm{\mu}}^{\top}\left(\alpha{\bm{1}}+JJ^{\top}{\bm{\mu}}\right)$
@@
Therefore, (20) and (21) are equivalent, with ${\bm{x}}=-J^{\top}{\bm{w}}$.
  Non-conflicting.
Observe that since in (21), $\alpha=0$ and ${\bm{x}}={\bm{0}}$ is feasible, the objective is non-positive and therefore $\alpha\varleq 0$. Substituting ${\bm{x}}=-J^{\top}{\bm{w}}$ in $J\cdot{\bm{x}}\preceq\alpha{\bm{1}}\preceq{\bm{0}}$ yields ${\bm{0}}\preceq JJ^{\top}{\bm{w}}$, i.e. ${\bm{0}}\preceq J\cdot\mathcal{A}_{\text{MGDA}}(J)$, so $\mathcal{A}_{\text{MGDA}}$ is non-conflicting.
  Linear under scaling.
With $J=\begin{bmatrix}2&0\\ 0&2\\ a&a\end{bmatrix}$, if $0\varleq a\varleq 1$, $\mathcal{A}_{\text{MGDA}}(J)=\begin{bmatrix}a\\ a\end{bmatrix}$. However, if $a\vargeq 1$, $\mathcal{A}_{\text{MGDA}}(J)=\begin{bmatrix}1\\ 1\end{bmatrix}$. This is not affine in $a$, so $\mathcal{A}_{\text{MGDA}}$ is not linear under scaling. In particular, if any row of $J$ is ${\bm{0}}$, $\mathcal{A}_{\text{MGDA}}(J)={\bm{0}}$. This implies that the optimization will stop whenever one objective has converged.
  Weighted.
By (19), $\mathcal{A}_{\text{MGDA}}$ is weighted.